OpenStudy (anonymous):

y=(x-4)(x+8)(8x-3) what are the y intercept and x intercept

OpenStudy (anonymous):

eqaute y with o to find x intercepts

OpenStudy (anonymous):

then equate x with o to find y intercepts

OpenStudy (anonymous):

Did that and came up with 0 and -3,-4 and 8 and the math system said the answers were wrong. Have no idea how to get the answer for this question

OpenStudy (jim766):

x-4=0 , x + 8 = 0 , 8x -3 = 0 solve each one to find the x intercepts

OpenStudy (anonymous):

x=4; -8 and 3/8 notas yours

OpenStudy (anonymous):

you got a mistake there

OpenStudy (anonymous):

Thank you but still really confused how do you get y?

OpenStudy (anonymous):

now use 0 instead of x

OpenStudy (anonymous):

96?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can find any graph by google, just type "graph for xxxx" in google search, where xxxx your equation without "y=". Just equation, in your case it is just (x-4)(x+8)(8x-3)

OpenStudy (anonymous):

Thank you so much I have another problem very similar and still cannot get this one correct either. Will try your suggestion above.

OpenStudy (anonymous):

Tried it and it did not work. Equation is x^3-1x^2-56x. Did not receive any search items that showed graphing all were solving a different equation. When I tried the substitute 0 for x and for y still getting wrong answer

OpenStudy (jim766):

x^3 - x^2 - 56x first find the GCF

OpenStudy (anonymous):

you got the one y intercept that is 0

OpenStudy (anonymous):

Thank you so much!! It worked now I have two problems left to find the polynomials. Is there a google site for that as well? Have been working on these 4 for almost a week with not getting a correct answer once. Polynomial problem is: degree 3 polynomial that goes through (5, -12.6) find the polynomial. Any ideas?

OpenStudy (anonymous):

this question was already solved

OpenStudy (anonymous):

No. This is all I was given and it states to write the polynomial. No idea how to do that?